# CS-Bench: A Comprehensive Benchmark for Large Language Models towards Computer Science Mastery

### v1

@@ 3.3 Qualitative Analysis
To investigate how the performance of models varies with the increase in parameter size, we examine several model families and plot the results in Figure 4 (a). It can be observed that although different families exhibit distinct performances, models within the same family consistently show improvement as the parameter size increases. However, as the model parameter size continues to increase, the performance gains from scaling diminish, resulting in diminishing returns in efficiency. For instance, the score in Qwen1.5 improves by 16.19% from 0.5B to 7B, by 7.11% from 14B to 72B, and by only 2.66% from 72B to 110B. Additionally, as shown in Figure 4 (b), when the parameter scale grows exponentially, the score approximately increases linearly. This indicates that in the CS field, the model’s performance also follows a logarithmic scale pattern. Given the substantial computational resources required for large-scale models, we aim to establish the relationship between model scales and scores to predict the performance of larger-scale models in the CS field by fitting smaller-scale model scores. Due to space limitations, the specific design and implementation of the fitting function are provided in Appendix E.2. Overall, we fit the functions of Llama2 and Qwen1.5 series based on models ranging from 7B to 70/72B. We validate the fitting function on Qwen-1.5 110B, where the predicted value (67.83%) closely matches the actual value (67.95%), enabling further predictions for theoretical models, even up to 1000B.
@@
Comparison between Zero-shot, Few-shot and COT Prompting.
To investigate the impact of few-shot prompts and chain of thought (COT [52]) on model performance, we evaluate model’s performance under 5-shot answer-only (AO) and 5-shot COT prompts in Figure 4 (c), where the prompt samples are sampled from the validation set and match the domain of the test questions. Given that model-generated results under 0-shot COT often don’t adhere to specific formats, making regular matching difficult, we omit 0-shot COT experiments, similar to C-Eval. Additionally, for Open-ended questions, since the answers include detailed explanations, 5-shot COT is the same as 5-shot AO. For all tested models, the 5-shot prompts show improvement compared to 0-shot, with average increases of 1.47% for 5-shot AO and 2.00% for 5-shot COT, respectively. Moreover, the efficacy of few-shot prompts in bringing improvements appears more pronounced in some robust models such as GPT-3.5 and GPT-4, owing to their superior in-context learning capabilities.
@@ 3.4 What’s the Relationship between CS, Math, and Code abilities of LLMs?
To explore the relationship between CS proficiency and the mathematical and coding capabilities of models, we investigate (1) the performance of general LLMs across the fields of Math, Code, and CS, and (2) the performance of LLMs specialized in Code and Math within the field of CS.
@@
In Figure 6, we illustrate how the models’ performance on CS-Bench varies with increasing scores on the Math datasets (GSM8K [53], MATH [35]) and Code datasets (HumanEval [54], MBPP [55]). We observe that the overall trend in CS-Bench performance closely aligns with changes in Math and Code scores, as indicated by a Pearson correlation coefficient [56] exceeding 0.9. Besides the general enhancement of diverse competencies that superior models typically bring, we consider this evidence to suggest a close correlation between CS proficiency and abilities in Math as well as Code.
Figure 6: The score changes on CS-Bench as LLM’s Math/Code score increases. $p$ denotes Pearson correlation coefficient. We obtain the scores on Math/Code datasets from [43].
@@
InternLM-Math-7B [57] Math 60.23 31.56 50.56 38.61 55.93 44.47 47.69 43.85 53.64 39.41 48.45
@@
DeepSeekMath-Instruct-7B [58] Math 63.98 34.82 55.13 39.64 61.26 42.16 45.29 42.69 56.68 39.67 50.49
@@
CodeLlama-7B [61] Code 58.90 36.15 45.46 36.24 52.87 26.23 44.35 25.33 50.36 31.09 43.34
Dolphcoder-7B [62] Code 50.13 36.47 34.71 34.36 41.78 23.92 40.03 28.35 41.40 30.82 37.54
WizardCoder-7B [63] Code 47.42 33.58 35.54 37.09 41.17 26.03 40.88 30.60 41.02 31.73 37.63
@@
CodeLlama-13B [61] Code 59.87 34.17 44.96 35.82 51.56 35.83 43.28 34.56 49.84 35.08 44.47
WizardCoder-13B [63] Code 50.80 32.98 38.69 35.27 43.42 28.34 40.88 34.29 43.27 32.59 39.38
@@
We present the results of the Math and Code expert LLMs in Tables 3 and 4. Compared to general Chat LLMs, expert LLMs usually sacrifice other abilities to boost proficiency in Math or Code, which is reflected in the lower overall performance of most expert LLMs. Therefore, we are more concerned with identifying the specific aspects of CS where Math and Code models show improvement. Regarding mathematics, InternLm-Math-7B improves InternLm2-7B’s performance in CO, CN, and OS reasoning tasks, while DeepseekMath exhibits significant improvements across all domains. According to [58], DeepseekMath effectively maintains general knowledge and reasoning ability during specialization. Conversely, MAammoTH and WizardMath perform poorly due to just fine-tuning on limited mathematical datasets, resulting in a significant decline in general knowledge and reasoning. The score changes in LLMs suggest that OS is most closely linked to mathematics, followed by CO, and lastly DSA and CN. In terms of Code, many Code models show significant improvements in DSA (especially knowledge) and OS (especially reasoning), such as CodeLlama and Dolphcoder. This indicates that the disciplines of DSA and OS are more closely related to code, thus enhancing knowledge and reasoning abilities in these directions, while CO and CN have lower relevance, leading to a decrease in scores. Finally, we observe that the enhancement brought about by small-scale expert LLMs compared to larger-scale LLMs is more pronounced (see CodeLlama-7B/13B, WizardCoder-7B/13B). We attribute this to the supplementary need for specific knowledge and reasoning capabilities in small-scale LLMs, whereas large-scale LLMs already encompass a greater breadth of knowledge and stronger reasoning abilities, resulting in diminishing gains from further training in specific domains.
@@ 4 Related Work
Given the powerful capabilities of LLMs, recent research has explored their potential applications across various industries and scientific fields, including finance [12], autonomous driving [10, 64], robotics [65, 9, 66], medicine [11, 67], and chemistry [8, 13]. Currently, studies exploring LLMs in the field of computer science fall into two main categories. The first category includes broad evaluation benchmarks covering various fields, such as MMLU [17], CMMLU [20], C-Eval [18], Xiezhi [21], and M3KE [19]. However, computer science constitutes only a small fraction of these benchmarks, accounting for less than 5% and lacking detailed CS-specific analysis. The second category focuses solely on exploring specific applications of LLMs within computer science, such as network topology [14], cybersecurity [68, 15], and software engineering [16, 69]. Nonetheless, there has been a persistent lack of comprehensive evaluation of LLMs’ foundational knowledge and reasoning abilities in computer science. To address this gap, we propose CS-Bench and conduct a thorough evaluation of LLMs, providing guidance for understanding and improving their performance in the CS field.
@@
Evaluating and understanding the capabilities of LLMs is a major focus within the NLP community. Researchers have extensively explored the capabilities of LLMs including planning [70], multilingual processing [71, 72], instruction following [73, 74], and cross-domain generalization [75, 76, 77]. Recently, there has been growing interest in LLMs’ abilities in mathematics [22, 23, 24, 25, 26, 27], code programming [63, 61, 62, 28, 29], and logical reasoning [30, 31, 32, 33]. While individual capabilities have been well-studied, research on their integrated application and interrelationships remains sparse. Different from [26], which investigates interactions between abilities during the supervised fine-tuning phase, we choose computer science as our research context. Given that computer science inherently integrates coding, mathematics, and reasoning, we utilize CS-Bench in this paper to deeply explore the relationship between LLMs‘ performance in computer science and their mathematical and coding abilities, aiming to advance cross-capability research and integrated analysis of LLM abilities.
@@ 5 Conclusion
In this work, we introduce CS-Bench, the first benchmark specifically designed to systematically analyze the knowledge and reasoning capabilities of mainstream LLMs in the field of computer science. Our evaluation of over 30 models highlights that even the top-performing GPT-4o has significant room for improvement in computer science. Further score-scale experiments and error type analyses provide directions for enhancing LLMs in the field. Moreover, our investigation into the relationship between computer science, mathematics, and coding demonstrates their close interconnections and provides valuable insights into LLMs’ cross-abilities and applications.
@@
Appendix A Limitations
@@
(2) Evaluation Limitations: In the CS-Bench evaluation experiments, we employ GPT-4 scoring to assess generative tasks such as fill-in-the-blank and open-ended tasks. This might lead to certain evaluation thresholds and costs. However, such issues only constitute about 20% of CS-Bench. Additionally, we provide an evaluation scheme that separates comprehension tasks from CS-Bench, allowing for automatic evaluations without the need for GPT-4.
(3) Language Limitations: CS-Bench are primarily focused on Chinese and English-dominated language environments, ensuring comprehensive and in-depth evaluations in these two language environments. However, for other non-Chinese and English language environments, its support and coverage are relatively weak, and further optimization and improvement are needed.
@@ Appendix B Broaden Impact
CS-Bench is anticipated to play a significant role in the field of computer science. LLMs, trained and evaluated with the aid of CS-Bench, can enhance the work efficiency of relevant professionals, enabling them to complete computer-related tasks, such as code review, error detection, and algorithm optimization, more quickly and accurately. Although this might result in the disappearance of some repetitive jobs, it could also create new career opportunities. In the realm of education, the CS-Bench dataset can serve as an effective teaching tool, assisting teachers in better explaining complex computer science concepts and techniques, and also enabling students to better understand and master this knowledge through practice.
@@ Appendix C More Details on CS-Bench
In C.1, we provide a detailed explanation of the design motivation and statistics for CS-Bench. In C.2, we present the distribution of question and answer lengths for each task in CS-Bench. In C.3, we provide a case example for each type under each dimension of CS-Bench.
C.1 Detailed Design Motivation and Statistics of CS-Bench

### v2

@@ 3.4 Qualitative Analysis
To investigate how model performance varies with parameter size, we examine several model families and plot the results in Figure 6 (a). Models within the same family consistently improve as parameter size increases, but performance gains diminish with larger models. For instance, the score in Qwen1.5 improves by 16.19% from 0.5B to 7B, by 7.11% from 14B to 72B, and by only 2.66% from 72B to 110B. Figure 6 (b) shows that when the parameter scale grows exponentially, the score increases approximately linearly, indicating a logarithmic scale pattern in the CS field. Given the substantial computational resources required for large-scale models, we aim to establish the relationship between model scales and scores to predict the performance of larger-scale models by fitting smaller-scale model scores. Due to space limitations, the specific design and implementation of the fitting function are provided in Appendix E.2. Overall, we fit the functions of Llama2 and Qwen1.5 series based on models ranging from 7B to 70/72B. We validate the fitting function on Qwen-1.5 110B, where the predicted value (67.83%) closely matches the actual value (67.95%), enabling further predictions for theoretical models, even up to 1000B.
Comparison between Zero-shot, Few-shot and COT Prompting.
To investigate the impact of few-shot prompts and chain of thought (COT (Wei et al., 2022)) on model performance, we evaluate model’s performance under 5-shot answer-only (AO) and 5-shot COT prompts in Figure 6 (c), where the prompt samples are sampled from the validation set and match the domain of the test questions. Given that model-generated results under 0-shot COT often don’t adhere to specific formats, making regular matching difficult, we omit 0-shot COT experiments, similar to C-Eval. Additionally, for Open-ended questions, since the answers include detailed explanations, 5-shot COT is the same as 5-shot AO. For all tested models, the 5-shot prompts show improvement compared to 0-shot, with average increases of 1.47% for 5-shot AO and 2.00% for 5-shot COT, respectively. Moreover, the efficacy of few-shot prompts in bringing improvements appears more pronounced in some robust models such as GPT-3.5 and GPT-4, owing to their superior in-context learning capabilities.
@@ 3.5 What’s the Relationship between CS, Math, and Code abilities of LLMs?
To explore the relationship between CS proficiency and the mathematical and coding capabilities of models, we investigate (1) the performance of general LLMs across the fields of Math, Code, and CS, and (2) the performance of LLMs specialized in Code and Math within the field of CS.  [^4] We leave the analysis from the model representation perspective in Appendix E.5.
@@
Figure 8: The score changes on CS-Bench as LLM’s Math/Code score increases. $p$ denotes Pearson correlation coefficient. We obtain the scores on Math/Code datasets from (Alibaba, 2024).
Figure 9: The consistency of model performance between CS-Bench and other benchmarks. Model scores from other benchmarks are sourced from their official papers/websites or the OpenLLM Leaderboard.
In Figure 9, we illustrate how the models’ performance on CS-Bench varies with increasing scores on the Math datasets (GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021b)) and Code datasets (HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021)). The overall trend in CS-Bench performance closely aligns with changes in Math and Code scores, as indicated by a Pearson correlation coefficient (Cohen et al., 2009) exceeding 0.9. Besides the general enhancement of diverse competencies that superior models typically bring, we consider this evidence suggests a strong correlation between CS proficiency and skills in Math and Code.
Another piece of evidence is the consistency between CS-Bench and other benchmarks, including non-CS scientific benchmarks (GPQA (Rein et al., 2023)), specialized benchmarks (Tool-use-eval (Alibaba, 2024), T-Eval (Chen et al., 2024), L-eval (An et al., 2023), BBH (Suzgun et al., 2022)), and general benchmarks (Chatbot Arena (Zheng et al., 2023), C-Eval(Huang et al., 2023b)). The results are shown in Figure 9. We observe that CS-Bench and agents, tool usage, and other fields have a low correlation (P$<$0.8), whereas their correlation with benchmarks involving long texts, multi-step reasoning, and general tasks is moderate (P$<$0.9). In contrast, the P-values for CS-Bench and benchmarks in mathematics/coding are all above 0.9, reaching as high as 0.96 in MBPP, which strongly supports the assertion that the model’s CS capabilities are closely related to its coding and mathematical abilities. Additionally, we find that Chatbot Arena, BBH, and L-eval include subsets related to mathematics and coding, resulting in their correlation with CS-Bench being higher than that of other non-Code/Math benchmarks.
@@
InternLM-Math-7B (Ying et al., 2024) Math 60.23 31.56 50.56 38.61 55.93 44.47 47.69 43.85 53.64 39.41 48.45
@@
DeepSeekMath-Instruct-7B (Shao et al., 2024) Math 63.98 34.82 55.13 39.64 61.26 42.16 45.29 42.69 56.68 39.67 50.49
@@
CodeLlama-7B (Rozière et al., 2024) Code 58.90 36.15 45.46 36.24 52.87 26.23 44.35 25.33 50.36 31.09 43.34
Dolphcoder-7B (Wang et al., 2024c) Code 50.13 36.47 34.71 34.36 41.78 23.92 40.03 28.35 41.40 30.82 37.54
WizardCoder-7B (Luo et al., 2023b) Code 47.42 33.58 35.54 37.09 41.17 26.03 40.88 30.60 41.02 31.73 37.63
@@
CodeLlama-13B (Rozière et al., 2024) Code 59.87 34.17 44.96 35.82 51.56 35.83 43.28 34.56 49.84 35.08 44.47
WizardCoder-13B (Luo et al., 2023b) Code 50.80 32.98 38.69 35.27 43.42 28.34 40.88 34.29 43.27 32.59 39.38
@@
We present the results of the Math and Code expert LLMs in Tables 3 and 4. Compared to general Chat LLMs, expert LLMs usually sacrifice other abilities to boost proficiency in Math or Code, which is reflected in the lower overall performance of most expert LLMs. Therefore, we are more concerned with identifying the specific aspects of CS where Math and Code models show improvement. Regarding mathematics, InternLm-Math-7B improves InternLm2-7B’s performance in CO, CN, and OS reasoning tasks, while DeepseekMath exhibits significant improvements across all domains. According to (Shao et al., 2024), DeepseekMath effectively maintains general knowledge and reasoning ability during specialization. Conversely, MAammoTH and WizardMath perform poorly due to just fine-tuning on limited mathematical datasets, resulting in a significant decline in general knowledge and reasoning. The score changes in LLMs suggest that OS is most closely linked to mathematics, followed by CO, and lastly DSA and CN.
In terms of Code, many Code models show significant improvements in DSA (especially knowledge) and OS (especially reasoning), such as CodeLlama and Dolphcoder. This indicates that the disciplines of DSA and OS are more closely related to code, thus enhancing knowledge and reasoning abilities in these directions, while CO and CN have lower relevance, leading to a decrease in scores. Finally, we observe that the enhancement brought about by small-scale expert LLMs compared to larger-scale LLMs is more pronounced (see CodeLlama-7B/13B, WizardCoder-7B/13B). We attribute this to the supplementary need for specific knowledge and reasoning capabilities in small-scale LLMs, whereas large-scale LLMs already encompass a greater breadth of knowledge and stronger reasoning abilities, resulting in diminishing gains from further training in specific domains.
@@ 4 Related Work
Given the powerful capabilities of LLMs, recent research has explored their potential applications across various industries and scientific fields, including finance (Zhao et al., 2024), autonomous driving (Huang et al., 2023a; Zhou et al., 2024), robotics (Yuan et al., 2023a; Xiao et al., 2023; Wang et al., 2024a), medicine (Zhou et al., 2023a; Vaid et al., 2024), and chemistry (Guo et al., 2023; Zhang et al., 2024a). Currently, studies exploring LLMs in the field of computer science fall into two main categories. The first category includes broad evaluation benchmarks covering various fields, such as MMLU (Hendrycks et al., 2021a), CMMLU (Li et al., 2024), C-Eval (Huang et al., 2023b), Xiezhi (Gu et al., 2024), and M3KE (Liu et al., 2023a). However, computer science constitutes only a small fraction of these benchmarks, accounting for less than 5% and lacking detailed CS-specific analysis. The second category focuses solely on exploring specific applications of LLMs within computer science, such as network topology (Donadel et al., 2024), cybersecurity (Ferrag et al., 2024; Murtuza, 2024), and software engineering (Marques et al., 2024; Dipongkor, 2024). Nonetheless, there has been a persistent lack of comprehensive evaluation of LLMs’ foundational knowledge and reasoning abilities in computer science. To address this gap, we propose CS-Bench and conduct a thorough evaluation of LLMs, providing guidance for understanding and improving their performance in the CS field.
@@
Evaluating and understanding the capabilities of LLMs is a major focus within the NLP community. Researchers have extensively explored the capabilities of LLMs including planning (Huang et al., 2024), multilingual processing (Lai et al., 2023; Bang et al., 2023), instruction following (Zhou et al., 2023b; Wang et al., 2023b), cross-domain generalization (Wang et al., 2023a; Song et al., 2023; Wang et al., 2024b) and safety (Mazeika et al., 2024; Diao et al., 2024; Mou et al., 2025). Recently, there has been growing interest in LLMs’ abilities in mathematics (Frieder et al., 2023; Collins et al., 2023; Wu et al., 2023; Yuan et al., 2023b; Dong et al., 2024; Liu et al., 2024), code programming (Luo et al., 2023b; Rozière et al., 2024; Wang et al., 2024c; Zhang et al., 2024b; Lin et al., 2024), and logical reasoning (Liu et al., 2023b; Saparov et al., 2023; Xu et al., 2023; Wu et al., 2024). While individual capabilities have been well-studied, research on their integrated application and interrelationships remains sparse. Different from (Dong et al., 2024), which investigates interactions between abilities during the supervised fine-tuning phase, we choose computer science as our research context. Given that computer science inherently integrates coding, mathematics, and reasoning, we utilize CS-Bench in this paper to deeply explore the relationship between LLMs’ performance in computer science and their mathematical and coding abilities, aiming to advance cross-capability research and integrated analysis of LLM abilities.
@@ 5 Conclusion
In this work, we introduce CS-Bench, the first benchmark specifically designed to systematically analyze the knowledge and reasoning capabilities of mainstream LLMs in the field of computer science. Our evaluation of over 30 models highlights that even the top-performing GPT-4o/OpenAI-o1 has significant room for improvement in computer science. Further score-scale experiments and error type analyses provide directions for enhancing LLMs in the field. Moreover, our investigation into the relationship between computer science, mathematics, and coding demonstrates their close interconnections and provides valuable insights into LLMs’ cross-abilities and applications.
Acknowledgements
This work was partially supported by:
  • The State Key Laboratory of Massive Personalized Customization System and Technology (No. H&C-MPC-2023-02-07(Q)).
  • The State Grid Technology Project (5700-202416236A-1-1-ZN) on “Research on Active Semantic Discovery Technology Based on SG-CIM and Its Application in Power Grid Equipment Supply Chain Optimization”.
  • China Unicom Software Research Institute under the “Framework Agreement for Seven Model Technology Research and Application Demonstration Projects (Software Development for Government Enterprise Content Generation) from 2024 to 2025” (No. 5500331818).
  • The National Natural Science Foundation of China (NSFC Nos. 62076031 and 62076036).
  • Zhongguancun Academy.
@@ Acknowledgements
Appendix A Limitations and Dataset Bias
@@
(2) Evaluation Limitations: In the CS-Bench evaluation experiments, we employ GPT-4 scoring to assess generative tasks such as fill-in-the-blank and open-ended tasks. This might lead to certain evaluation thresholds and costs. However, such issues only constitute about 20% of CS-Bench. Additionally, we provide an evaluation scheme that separates comprehension tasks from CS-Bench, allowing for automatic evaluations without the need for GPT-4. We also explore the effect of different scoring models and find that even slightly inferior models produce score rankings consistent with GPT-4, while significantly reducing costs.
(3) Language Limitations: CS-Bench are primarily focused on Chinese, French, German, and English language environments, ensuring comprehensive and in-depth evaluations in these language environments. However, for other environments, its support and coverage are relatively weak, and further optimization and improvement are needed.
The main biases of CS-Bench can be categorized into two aspects:
(1) Difficulty Level: Overall, the benchmark reflects a university-level difficulty.
(2) Task Focus: It emphasizes knowledge- and reasoning-based questions rather than specific real-world production scenarios in computer science.
The impact of difficulty on LLM evaluation is: lower difficulty tends to narrow the score differences between models, while higher difficulty amplifies these differences. It is worth noting that although CS-Bench maintains an overall university-level difficulty, the diversity of difficulty across questions remains significant, ranging from simple definitions in data structures to challenging computer network application problems. This is reflected in the model score range of 39.86% to 72.29%, demonstrating CS-Bench’s effectiveness in distinguishing LLMs’ capabilities in computer science and presenting challenges even for the best-performing models.
Regarding task scenarios, a key motivation for CS-Bench is that “some prior work has explored LLMs in specific applications such as cybersecurity and software engineering.” However, as stated in Section 4, there has been a lack of comprehensive evaluations of LLMs’ fundamental knowledge and reasoning capabilities in computer science. Besides, considering the extensive research already conducted by the LLM community on coding abilities, we view programming as a separate focus area. Therefore, CS-Bench does not include programming questions but instead focuses on the relationship between CS capabilities and coding abilities.
@@ Appendix B Broaden Impact
CS-Bench is anticipated to play a significant role in the field of computer science. LLMs, trained and evaluated with the aid of CS-Bench, can enhance the work efficiency of relevant professionals, enabling them to complete computer-related tasks, such as code review, error detection, and algorithm optimization, more quickly and accurately. Although this might result in the disappearance of some repetitive jobs, it could also create new career opportunities. In the realm of education, the CS-Bench dataset can serve as an effective teaching tool, assisting teachers in better explaining complex computer science concepts and techniques, and also enabling students to better understand and master this knowledge through practice. However, we should also be cautious of potential risks associated with CS-Bench, such as exam cheating by students in the CS field, which requires management mechanisms and logging to prevent misuse.
@@ Appendix C More Details on CS-Bench
In C.1, we introduce the details of data collection and processing for CS-Bench. In C.2, we provide a detailed explanation of the design motivation and statistics for CS-Bench. In C.3, we present the distribution of question and answer lengths for each task in CS-Bench. In C.4, we provide a case example for each type under each dimension of CS-Bench.
C.1 Details of Data Collection and Processing.
Our data mainly comes from three sources: 1. Public channels providing computer science-related questions, such as professional exams and practice tests555e.g., https://github.com/CodePanda66/CSPostgraduate-408, https://github.com/ddy-ddy/cs-408. 2. Knowledge-based questions manually extracted and adapted by professionals from various academic-permitted CS-related blog posts 666e.g., https://www.wikipedia.org/, https://www.cnblogs.com/, https://www.csdn.net/,
https://zhuanlan.zhihu.com
. 3. Questions constructed from non-public teaching materials and exam papers authorized by the author’s affiliated institution. For resources sourced from the internet, we manually extract the knowledge points and questions. For physical materials, we use Optical Character Recognition (OCR) to obtain the data. We then ensure the accuracy of the collected data through manual cross-checking. The data collection was carried out by a team of five students with bachelor’s degrees in computer science. We provided each person with adequate wages, significantly above the local minimum wage standard. Based on whether the questions require in-depth reasoning and computation, we label each question as either knowledge-type or reasoning-type. Additionally, we tag each instance with domain and task type labels. For English data, we used GPT-4 to translate the Chinese instances into English, French, and German, followed by manual verification.
C.2 Detailed Design Motivation and Statistics of CS-Bench